OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f ° g)(-5). A. 23 B. –53 C. –9 D. 3

OpenStudy (anonymous):

open circle means apply g(-5) first, and then use the result for the f(x) function

OpenStudy (anonymous):

do i use -5 for f(x) as well?

OpenStudy (anonymous):

no, you use the result of g(-5) as x in f(x)

OpenStudy (anonymous):

kind of like f(g(-5))

OpenStudy (anonymous):

so g(-5) = -4*(-5) +3 = 23 f(23) = -2*(23) -7 = -53 so your answer is -53

OpenStudy (anonymous):

gracias mi amigo